# The Limits of Price Discrimination Under Privacy Constraints

### v3

@@ 3.1 Illustration of the set $\mathcal{S}$
Figures 3(b) and 3(c) correspond to the case $\alpha^{*}\geq\eta\geq 1/2$. Notice that $\eta\geq 1/2$ implies $t^{*}\geq\eta$, and hence in this case, we could have either $\alpha^{*}\geq t^{*}$ or $\alpha^{*}\leq t^{*}$, depending on the value of $\beta$. Let $\tilde{\beta}$ be the value of $\beta$ for which $t^{*}=\alpha^{*}$, i.e.,
@@ 3.2 Does imposing privacy hurt or help the producer and the consumers?
We next discuss that imposing privacy always hurts the producer but, interestingly, may hurt or help consumers. We do so by comparing the set $\mathcal{S}$ with the non-private case in Figures 3(d) and 4(d), studied in Bergemann et al. [2015].
@@ 3.3 The indirect effect of privacy
With $\eta\leq 1/2$, the producer is more inclined to set the market price at $v_{2}$. However, the assumption $\alpha^{*}\leq\eta$ indicates a scarcity of high-value customers, making $v_{1}$ the optimal uniform price. This is why in Figures 4(a) and 4(b), we observe points below the uniform pricing line $\tilde{T}\tilde{R}$. This effect is particularly intensified when $\beta$ is high, i.e., Figure 4(a), as the threshold for setting the price high (i.e., $t^{*}$) approaches zero. This adversely impacts consumers since consumers gain utility only when they have a high value and the marker is priced at the low value $v_{1}$.
@@ 4 Price discrimination limits under privacy: the general case
Consider the case $K=5$ with values $[v_{k}]_{k=1}^{5}=[0.8,2,3,4.2,5]$ and $\beta=0.3$. Figure 6 illustrates $\mathcal{S^{\prime}}$ for three distinct aggregated market values $x^{*}$. Specifically, the aggregated markets associated with Figures 5(a), 5(b), and 5(c) are $(0.2,0.1,0.4,0.2,0.1)$, $(0.2,0.3,0.2,0.2,0.1)$, and $(0.2,0.1,0.1,0.05,0.55)$, respectively. The dashed triangle $TQR$ in the figures represents the non-private case, in which, as previously stated, the set $\mathcal{S}^{\prime}$ coincides with the set $\mathcal{S}$.
@@
We defer the proof of this proposition to the appendix. Revisiting Example 1, we find that while the condition (30) is not met for $x^{*}$ corresponding to Figure 5(a), it is fulfilled for the aggregated markets associated with Figures 5(b) and 5(c). Consequently, the set $\mathcal{S}^{\prime}$ intersects with line $TR$ in the former case without crossing it, whereas in the latter two cases, it extends below this line.
For the case $K=2$, the condition (30) translates into either $\alpha^{*}\geq\eta\geq 1/2$ or $\alpha^{*}\leq\eta\leq 1/2$. The former condition, $\alpha^{*}\geq\eta\geq 1/2$, corresponds to Figures 3(b) and 3(c). It is important to note that these figures depict the set $\mathcal{S}$, which is derived from the set $\mathcal{S}^{\prime}$ after applying the scaling and shift described in Proposition 3. Specifically, line $TR$ in $\mathcal{S}^{\prime}$ corresponds to line $\tilde{T}\tilde{R}$ in these figures. Consequently, the lowest point of $\mathcal{S}^{\prime}$ maps to point $C$ that is located below the line $\tilde{T}\tilde{R}$ in both cases. In contrast, for Figure 3(a), where the condition (30) is not met, the lowest point of $\mathcal{S}^{\prime}$ maps to point $A$, which lies precisely on the line $\tilde{T}\tilde{R}$. Similarly, for the case $\alpha^{*}\leq\eta\leq 1/2$, as illustrated in Figures 4(a) and 4(b), the lowest point of $\mathcal{S}^{\prime}$ maps to point $A$ which is situated below the line $\tilde{T}\tilde{R}$.
@@ A.1 Proof of Proposition 1
A.2 Proof of Lemma 1
Notice that the producer’s utility by setting the price equal to $v_{1}$ is given by $\mathcal{U}_{p}(v_{1}|\hat{x})=v_{1}$ as all the consumer will buy the product at this price. For price $v_{2}$, the expected utility is given by
@@ A.3 Proof of Proposition 2
Consider a market $x=(1-\alpha,\alpha)$ and a pricing rule ϕ()italic-$\phi(\cdot)$ that sets the price of a market $v_{2}$ with probability $\delta$ in the case of observing $\hat{\alpha}=t^{*}$.
@@
[…]
@@
Now we establish the converse. Suppose $\alpha_{1},\alpha_{2},\gamma_{1}$, and $\gamma_{2}$ are given, satisfying the conditions on the right-hand side of (35). We will show that there exists a segmentation $\sigma\in\Sigma$ and an optimal pricing rule $\phi_{\delta}(\cdot)$ such that (36) is satisfied. If $\alpha_{1}<t^{*}$ and $\alpha_{2}>t^{*}$, we can consider a segmentation with two markets $(1-\alpha_{1},\alpha_{1})$ and $(1-\alpha_{2},\alpha_{2})$, assigned probabilities $\gamma_{1}$ and $\gamma_{2}$ respectively, and apply any optimal pricing rule to satisfy (36). If $\alpha_{1}=t^{*}$ and $\alpha_{2}>t^{*}$, then $\phi_{0}(\cdot)$ should be taken as the optimal pricing rule. Conversely, if $\alpha_{1}<t^{*}$ and $\alpha_{2}=t^{*}$, $\phi_{1}(\cdot)$ becomes the optimal pricing rule. The only remaining scenario is $**\alpha_{1}=\alpha_{2}=t^{*}*$, where we take $\phi_{\delta}(\cdot)$ with $\delta=\gamma_{2}$ as the optimal pricing rule. This completes the proof. ∎
@@ A.12 Proof of Proposition 5
where $\bm{z}^{*}_{\beta}$ is the solution of (A.12). Next, using the Karush–Kuhn–Tucker conditions Bertsekas [1997], we know that (i) $\lambda_{i,j}\geq 0$, and (ii) $\lambda_{i,j}>0$ if and only if the condition (72c) is active, i.e.,

### v4

@@ 3.1 Illustration of the set $\mathcal{S}$
Figures 4(b) and 4(c) correspond to the case $\alpha^{*}\geq\eta\geq 1/2$. Notice that $\eta\geq 1/2$ implies $t^{*}\geq\eta$, and hence in this case, we could have either $\alpha^{*}\geq t^{*}$ or $\alpha^{*}\leq t^{*}$, depending on the value of $\beta$. Let $\tilde{\beta}$ be the value of $\beta$ for which $t^{*}=\alpha^{*}$, i.e.,
@@ 3.2 Does imposing privacy hurt or help the producer and the consumers?
We next discuss that imposing privacy always hurts the producer but, interestingly, may hurt or help consumers. We do so by comparing the set $\mathcal{S}$ with the non-private case in Figures 4(d) and 5(d), studied in Bergemann et al. [2015].
@@ 3.3 The indirect effect of privacy
With $\eta\leq 1/2$, the producer is more inclined to set the market price at $v_{2}$. However, the assumption $\alpha^{*}\leq\eta$ indicates a scarcity of high-value customers, making $v_{1}$ the optimal uniform price. This is why in Figures 5(a) and 5(b), we observe points below the uniform pricing line $\tilde{T}\tilde{R}$. This effect is particularly intensified when $\beta$ is high, i.e., Figure 5(a), as the threshold for setting the price high (i.e., $t^{*}$) approaches zero. This adversely impacts consumers since consumers gain utility only when they have a high value and the marker is priced at the low value $v_{1}$.
@@ 4 Price discrimination limits under privacy: the general case
Consider the case $K=5$ with values $[v_{k}]_{k=1}^{5}=[0.8,2,3,4.2,5]$ and $\beta=0.3$. Figure 6 illustrates $\mathcal{S^{\prime}}$ for three distinct aggregated market values $x^{*}$. Specifically, the aggregated markets associated with Figures 6(a), 6(b), and 6(c) are $(0.2,0.1,0.4,0.2,0.1)$, $(0.2,0.3,0.2,0.2,0.1)$, and $(0.2,0.1,0.1,0.05,0.55)$, respectively. The dashed triangle $TQR$ in the figures represents the non-private case, in which, as previously stated, the set $\mathcal{S}^{\prime}$ coincides with the set $\mathcal{S}$.
@@
We defer the proof of this proposition to the appendix. Revisiting Example 1, we find that while the condition (31) is not met for $x^{*}$ corresponding to Figure 6(a), it is fulfilled for the aggregated markets associated with Figures 6(b) and 6(c). Consequently, the set $\mathcal{S}^{\prime}$ intersects with line $TR$ in the former case without crossing it, whereas in the latter two cases, it extends below this line.
For the case $K=2$, the condition (31) translates into either $\alpha^{*}\geq\eta\geq 1/2$ or $\alpha^{*}\leq\eta\leq 1/2$. The former condition, $\alpha^{*}\geq\eta\geq 1/2$, corresponds to Figures 4(b) and 4(c). It is important to note that these figures depict the set $\mathcal{S}$, which is derived from the set $\mathcal{S}^{\prime}$ after applying the scaling and shift described in Proposition 4. Specifically, line $TR$ in $\mathcal{S}^{\prime}$ corresponds to line $\tilde{T}\tilde{R}$ in these figures. Consequently, the lowest point of $\mathcal{S}^{\prime}$ maps to point $C$ that is located below the line $\tilde{T}\tilde{R}$ in both cases. In contrast, for Figure 4(a), where the condition (31) is not met, the lowest point of $\mathcal{S}^{\prime}$ maps to point $A$, which lies precisely on the line $\tilde{T}\tilde{R}$. Similarly, for the case $\alpha^{*}\leq\eta\leq 1/2$, as illustrated in Figures 5(a) and 5(b), the lowest point of $\mathcal{S}^{\prime}$ maps to point $A$ which is situated below the line $\tilde{T}\tilde{R}$.
@@ A.1 Proof of Proposition 1
A.2 Connections with differential privacy
In the context of pricing based on histograms or distribution of values, the dataset is the set of consumers’ values, and the algorithm’s output is the selected market price. Thus, the goal of this privacy framework is to ensure that one cannot infer too much about the consumers’ values by observing the optimal price. Note that this addresses a different aspect of privacy compared to that which we discussed previously. In our setting, the main concern is not the producer itself but rather what an adversary can infer about the consumers’ values by observing the price set for them. For instance, if one observes that the insurance premium is set high for a zip code, one may infer that residents of that zip code often require medical assistance. The application of differential privacy to pricing has also been studied in the context of revenue management [see, e.g., Lei et al., 2023]. A caveat is that differential privacy is typically applicable to settings with a finite number of users, as opposed to our setting with a continuum of consumers. We discuss this matter in Section A.2, provide an adjusted definition of differential privacy, and demonstrate how our privacy mechanism (3) is, in fact, differentially private.
Let us first recall the definition of differential privacy for an algorithm $\mathcal{A}$ over a set of $n$ users’ data sets, denoted by $\bm{z}=\{z_{1},\cdots,z_{n}\}$ [see Dwork et al., 2014, for a detailed discussion of the definition and its applications].
Definition 3.
A randomized algorithm $\mathcal{A}:\mathcal{Z}^{n}\to\mathcal{Y}$ is called $\varepsilon$-differentially private if, for any measurable set $\mathcal{O}$, we have
${\mathbb{P}(\mathcal{A}(\bm{z})\in\mathcal{{O}})}\leq e^{\varepsilon}~{}{\mathbb{P}(\mathcal{A}(\bm{z^{\prime}})\in\mathcal{{O}})},$ (35)
for any two neighboring datasets $\bm{z}$ and $\bm{z^{\prime}}$ that only differ in one coordinate.
Here, $\varepsilon$ quantifies the privacy loss. The higher $\varepsilon$ is, it means the output is more sensitive to the user’s data, meaning that it provides a weaker privacy guarantee.
One challenge here is that, in our case, we work with a continuum of consumers, and hence, changing one user’s data is not meaningful. A natural proposal is to consider two markets that are close enough as a surrogate for neighboring distributions. In particular, we can use a distribution distance function $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{\geq 0}$ as a measure of the closeness of markets and say two markets $x$ and $x^{\prime}$ are neighboring markets if $d(x,x^{\prime})\leq\Delta$ for some choice of $\Delta>0$. We, therefore, use the following adjusted definition. Recall that $\phi(\cdot)$ denotes the optimal pricing rule and $\mathcal{X}$ represents the set of markets.
Definition 4.
A randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{X}$ is called $\varepsilon$-differentially private if, for any value $v_{i}\in\mathcal{V}$, we have
$\mathbb{P}(\phi(\mathcal{M}(x))=v_{i})\leq e^{\varepsilon}~{}\mathbb{P}(\phi(\mathcal{M}(x^{\prime}))=v_{i}),$ (36)
for two markets $d(x,x^{\prime})\leq\Delta$.
Now, let us see how our mechanism fits into this definition.
Proposition 7.
The mechanism $\mathcal{M}_{\beta}(\cdot)$ is $\varepsilon$-differentially private with
$\varepsilon=\max_{i:\text{Vol}(X_{i}^{\beta})\neq 0}\frac{1-\beta+\beta\text{Vol}(X_{i}^{\beta})}{\beta\text{Vol}(X_{i}^{\beta})},$ (37)
where $X_{i}^{\beta}$ is defined in (19).
Proof.
To simplify the notation, let us denote $\mathcal{M}_{\beta}(x)$ by $\hat{x}$. Then, we have
$\mathbb{P}(\phi(\hat{x})=v_{i})=\mathbb{P}(\hat{x}\in X_{i}^{\beta})=(1-\beta)\mathbbm{1}(x\in X_{i}^{\beta})+\beta\text{Vol}(X_{i}^{\beta}).$
By changing $x$, the term $\mathbbm{1}(x\in X_{i}^{\beta})$ could change from zero to one, but the second term $\beta\text{Vol}(X_{i}^{\beta})$ remains unchanged. This, consequently, implies the result. ∎
A.3 Proof of Proposition 2
Let $\nu_{\beta}(x,\hat{x})$ denote the joint distribution of $x$ and $\hat{x}$. With a slight abuse of notation, we denote the distribution of $\hat{x}$ condition on $x$ by $\nu_{\beta}(\hat{x}|x)$. Note that the term in (5) that we wish to maximize is equal to
[…]
Note that, given the privacy mechanism $\mathcal{M}_{\beta}(\cdot)$, $\int_{\mathcal{X}}\nu_{\beta}(\hat{x}|x)\log(\nu_{\beta}(\hat{x}|x))$ is independent of $x$, and therefore, the first term in (41) does not depend on the choice of $\pi(\cdot)$. The second term is the entropy of $\mu_{\beta}(\hat{x})$ and we know that it is maximized when $\mu_{\beta}(\cdot)$ is the uniform distribution, which is the case when $\pi(\cdot)$ is the uniform distribution. This completes the proof.
A.4 Proof of Lemma 1
Notice that the producer’s utility by setting the price equal to $v_{1}$ is given by $\mathcal{U}_{p}(v_{1}|\hat{x})=v_{1}$ as all the consumers will buy the product at this price. For price $v_{2}$, the expected utility is given by
@@ A.5 Proof of Proposition 3
Consider a market $x=(1-\alpha,\alpha)$ and a pricing rule $\phi(\cdot)$ that sets the price of a market $v_{2}$ with probability $\delta$ in the case of observing $\hat{\alpha}=t^{*}$.
@@
$\qquad\alpha_{1}\in[0,\min\{t^{*},\alpha^{*}\}],\alpha_{2}\in[\max\{t^{*},\alpha^{*}\},1],\gamma_{1}\text{ and }\gamma_{2}\in[0,1],\gamma_{1}+\gamma_{2}=1,\gamma_{1}\alpha_{1}+\gamma_{2}\alpha_{2}=\alpha^{*}\biggr{\}}$
@@
Now we establish the converse. Suppose $\alpha_{1},\alpha_{2},\gamma_{1}$, and $\gamma_{2}$ are given, satisfying the conditions on the right-hand side of (43). We will show that there exists a segmentation $\sigma\in\Sigma$ and an optimal pricing rule $\phi_{\delta}(\cdot)$ such that (44) is satisfied. If $\alpha_{1}<t^{*}$ and $\alpha_{2}>t^{*}$, we can consider a segmentation with two markets $(1-\alpha_{1},\alpha_{1})$ and $(1-\alpha_{2},\alpha_{2})$, assigned probabilities $\gamma_{1}$ and $\gamma_{2}$ respectively, and apply any optimal pricing rule to satisfy (44). If $\alpha_{1}=t^{*}$ and $\alpha_{2}>t^{*}$, then $\phi_{0}(\cdot)$ should be taken as the optimal pricing rule. Conversely, if $\alpha_{1}<t^{*}$ and $\alpha_{2}=t^{*}$, $\phi_{1}(\cdot)$ becomes the optimal pricing rule. The only remaining scenario is $\alpha_{1}=\alpha_{2}=t^{*}*$, where we take $\phi_{\delta}(\cdot)$ with $\delta=\gamma_{2}$ as the optimal pricing rule. This completes the proof. ∎
@@ A.14 Proof of Proposition 6
where $\bm{z}^{*}_{\beta}$ is the solution of (80). Next, using the Karush–Kuhn–Tucker conditions Bertsekas [1997], we know that (i) $\lambda_{i,j}\geq 0$, and (ii) $\lambda_{i,j}>0$ if and only if the condition (80c) is active, i.e.,